OpenStudy (kbase4ek):

Please help, I'll fan and medal Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (kbase4ek):

@mathmate

OpenStudy (kbase4ek):

@mathstudent55 ???

OpenStudy (mathmate):

Hint: Where there is no limitations on the parameters, r can be positive or negative, and \(\theta\) can take on positive and negative values. However, if \(0\le \theta \lt 360\), then \(\theta\) is unique, and you're left with only one option.

OpenStudy (kbase4ek):

Choices: (2 square root of 2, 225°), (-2 square root of 2, 45°) (2 square root of 2, 135°), (-2 square root of 2, 315°) (2 square root of 2, 315°), (-2 square root of 2, 135°) (2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (kbase4ek):

@mathmate

OpenStudy (mathmate):

If you need more explanations on the hints, I'll be glad to help.